OpenStudy (anonymous):

pls help. Asley and amy are seling pies. There are apple and lemon. Ashley sold 12 apple and 10 lemon pies for a total of $174. Amy sold 11 apple and 5 lemon pies for a total of $122. What is the individual cost for the apple and the cost for the lemon pie? WILL GIVE MEDAL

OpenStudy (triciaal):

A = apple L =lemon 12A + 10L = 174 11A + 5L = 122 option 1 multiply the 2nd eqn by 2 and subtract from the first (12-22)A = 174 -2(122) A = 7 substitute for A in either eqn in eqn 2; 5L = 122 - 77 L = 45/5 = 9 check in eqn 1 12 x 7 + 10 x 9 = 84 + 90 = 174